Circular No. 6086

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

X-RAY NOVA IN SCORPIUS R. M. Hjellming and M. P. Rupen, National Radio Astronomy Observa-

tory, report: “We have obtained 13 epochs of VLA 22.5-GHz images (Aug. 18-Sept. 29) of the resolved radio source identified with GRO J1655−40. The earliest maps show a double source evolving to a triple; this, together with VLBA images at much higher resolution taken on Aug. 18, 22 and 25 at 1.7 GHz, supports the conclusion that we are seeing twin-jet ejection of relativistic plasma from the central component. The originally-reported (IAUC 6073) total separation rate of 0′′.056 per day can now be separated into the oppositely-directed linear motions of the two outer components, the one to the northeast moving at 0′′.040 ± 0′′.005 per day, the one to the southwest at 0′′.020 ± 0′′.005 per day. An additional (fourth) component appeared northeast of the core in our maps on Sept. 11, our first observation following the radio outburst on Sept. 6 (IAUC 6077); this component also appears to be moving at 0′′.040 per day, along the same trajectory as the original northeast blob. Assuming a distance of 3.5 kpc (IAUC 6062), these expansion rates indicate apparent transverse speeds with respect to the central component of 0.8c for the northeast components and 0.4c for the one to the southwest. This relativistic motion is supported by the relative brightnesses of the components, with the blobs to the northeast being more prominent than those to the southwest. Taking the apparent difference in speeds as a relativistic effect, we find an intrinsic speed for the ejecta of (0.33 ± 0.12)c/ cos θ, where θ is the inclination of the trajectory to the line-of-sight. This implies that the angle to the line-of-sight is 70^{o} or less, suggesting there should be a significant radial-velocity component in any spectral lines associated with the jet. The preliminary VLBA images show that all three of the components seen at early times had complex structure elongated along the apparent directions of motion. The central component continues to show variations in flux density (based on multifrequency VLA monitoring) that are most prominent at frequencies above 8 GHz. Given the high apparent velocities, frequent optical spectroscopy is urgently needed.”

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Total magnitude estimates: Sept. 3.04 UT, 9.0 (J. D. Shanklin, Cam-

bridge, England, 0.20-m refractor); 6.22, 8.2 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 9.20, 8.3 (Spratt); 12.20, 8.8 (Spratt); 25.47, 12.8 (A. Nakamura, Kuma, Shikoku, Japan, 0.60-m reflector + unfiltered CCD); 27.46, 10: (P. Camilleri, Cobram, Victoria, Australia, 20 × 80 binoculars).

1994 October 1 Brian G. Marsden